# Interaction-Augmented Instruction: Modeling the Synergy of Prompts and Interactions in Human-GenAI Collaboration

### v1

@@ 4.3. Cross-paradigm Insights and Suggestions
Taken together, the twelve atomic paradigms show how interaction timing and artifact availability jointly govern how user intent is expressed. Below we distill concrete design considerations for UI designers and GenAI researchers.
@@
C1. Considering the timing for interaction and GenAI invocation based on intent clarity. Interaction timing indexes whether a task’s intent is knowable up front (pre-invocation) or emerges through exploration (post-invocation). Pre-invocation paradigms assume the user can articulate explicit and clear intent in advance for generating or manipulating artifacts, such as clear coding structure and logic (Feng et al., 2024), or explicit image editing areas and tasks  (Liu et al., 2025). Post-invocation paradigms assume underspecified or exploratory goals: the system first initiates interactions to further clarify user intents (Wang et al., 2024a) or produces outputs for steering and refinement (Zhang et al., 2025). Designers should align timing with the task’s nature: employ pre-invocation paradigms for tasks requiring precision and auditable outcomes, but prioritize lightweight post-invocation paradigms when the goal is creative ideation and iterative refinement of an ambiguous intent.
@@
C2. Grounding interactions with artifacts for referential intent. Instructions in natural language alone can be ambiguous (e.g., make the flower brighter” without specifying which flower). The presence of an artifact shifts intent expression from purely linguistic descriptions to concrete, grounded references. By enabling users to select, highlight, or annotate artifact fragments, systems can reduce ambiguity and provide a stable foundation for commands (Singh et al., 2024; Chen et al., 2023). The consideration also aligns with design principles for AI-instruments by Riche et al.  (Riche et al., 2025), where they proposed that text prompts should be grounded in other artifacts.
@@
C3. Expressing and materializing intent with an appropriate form. In these design paradigms, a key design choice is how user intent is materialized to augment the original text prompt. The first way is to directly edit or extend the text prompt with interactions (Table 2-P1 and P5). An example is DreamSheets (Almeda et al., 2024), where users can combine prompt fragments easily with interactions. It provides a simple and direct way to improve original prompts but lacks the power to cater complex intent, such as non-linguistic or referential intent. To handle more complex intent, an approach is to leverage interactions to introduce additional non-linguistic information to text and form augmented instructions, such as structural or parametric information (Table 2-P2, P3, P6, P7, P9 and P11). For example, CoPrompt (Feng et al., 2024) allows users to drag and drop prompt fragments to form a multi-level list of instructions. It materializes the non-linguistic intent expression in a convenient way with suitable interaction design. Regarding the referential intent, the key interaction is to link text prompts with artifacts and generate augmented instructions (Table 2-P4, P10). It addresses the challenges in describing the linkage between texts and artifacts by materializing users’ intent with direct interactions on artifacts. Lastly, a unique case is to generate new interaction widget by text prompts and facilitate intent expression through direct manipulation (Table 2-P8, P12). GenAI creates reusable and persistent interactive widgets to materialize the meta intent and allows follow-up similar intent expression with simple interactions with widgets. A notable example is DynaVis. The meta intent like changing visual element colors can be materialized as a widget. The specific color to use can be directly selected by interactions. These different combinations of interactions and instructions provide a huge space to suit different intent expression and materialization. Beyond timing and artifacts availability, they serves as a key consideration for designers.
@@
C4. Reusing, chaining, and innovating atomic design paradigms for adapting interaction design to new scenarios. Our summarized paradigms and the considerations C1-C3 above provide a concise mapping from task requirements to interface affordances. Designers can start from intent clarity and artifact availability to select appropriate paradigms for interface implementation. For example, they can use the pre-invocation, artifact-grounded paradigm (P4) for tasks that start from existing artifacts and demand precision. They can apply post-invocation, prompt-driven paradigms (P5-P8) for ideation and exploration. The should also consider the intent type as C3 mentions. Crucially, the twelve paradigms are not mutually exclusive: they can be flexibly combined and chained within a single application, enabling systems to shift fluidly between scaffolding, refinement, and repurposing workflows (Gmeiner et al., 2025; Wang et al., 2024a). Effective interfaces thus treat paradigms as composable building blocks rather than rigid templates, supporting diverse user needs. For example, IntentTagger (Gmeiner et al., 2025) introduces small, atomic intent tags enabling micro-prompting and region-level edits, illustrating chaining and covering multiple paradigms (e.g., P2, P4, P5, P7, P9). Reusing and chaining paradigms are not the end. Comparing the summarized paradigms and the entire IAI model, we can notice there are plenty of potential atomic paradigms that have not been explored. In future, interaction designers may start from existing paradigms to designing new ones for innovation.
@@ 5. Usage Scenario
The generative power of our proposed interaction-augmented instruction model and the twelve summarized atomic paradigm graphs extend beyond theoretical formulation and provides a practical design framework for human–GenAI communication (Beaudouin-Lafon et al., 2021). This framework can guide both the iterative improvement of existing tools and the creation of novel interfaces tailored to emerging user scenarios. To illustrate the model’s applicability and impact, we present four usage scenarios to highlight how our approach can inspire, structure, and accelerate design decisions in real-world contexts, bridging the gap between conceptual models and actionable interface innovation.
@@ 5.1. Usage Scenario 1: Extending Pipelines through Chained Paradigm Graphs
One way our model supports innovation is by extending existing human-GenAI interfaces through the chaining of additional atomic paradigm graphs onto their current workflows (Sec. 4.3-C4). This perspective treats current system interfaces not as static endpoints but as expandable foundations, where other paradigms can be strategically layered to address ambiguity, improve alignment, and ultimately enable more effective human-GenAI collaboration.
For example, when designing an ideal tool for data analysts to generate and edit visualizations, natural language can ease the burden of translating design requirements into initial visualizations, while graphical interactions can make subsequent adjustments and fine-grained editing more efficient. DynaVis (Vaithilingam et al., 2024) (Fig. 4-P8) illustrates this synergy by augmenting natural language interfaces with dynamic widgets that support iterative refinement. Yet, at the same time, early-stage ambiguity in user intent often remains a challenge. Fig. 6 illustrates this, imagine a data analyst at a movie company exploring market trends for the second quarter across U.S. cinemas. The analyst might ask: “Show me a trend of the most popular movie in the recent three months.” Directly generating data insights to such fuzzy user questions (e.g., “the most popular”) may not fully align with the user’s intent, frequently leading to time-consuming post-generation refinements. Using the IAI model and the atomic paradigm graphs, system designers can find and introduce a pre-generation disambiguation paradigm (P5, AI-driven Prompt Suggestion) before P8. Instead of immediately generating a visualization, the system could first initiate a clarification step, asking the analyst what “popular” should mean in this context, by ticket sales, IMDB rating, box office revenue, or number of screenings. This step structures the workflow to clarify ambiguous prompts upfront, improving alignment, reducing back-and-forth iterations, and making subsequent widget-driven refinement more precise.
@@ 5.2. Usage Scenario 2: Refining Paradigms by Adjusting Graph Structures
In addition to strengthening workflows at a macro level by chaining different paradigms, the flexibility of the relations among entities in the IAI model also enables micro-level edits within a single paradigm. Such refinements can be realized by adjusting the relations among entities in existing paradigms to improve precision and control.
@@ 5.3. Usage Scenario 3: Applying the IAI model to Emerging Scenarios
Consider Data Director (Shen et al., 2024a), a multi-agent system that automatically generates animated data videos from data tables with agent roles such as Data Analyst and Designer (Fig. 8). Although such end-to-end automation can rapidly deliver initial outputs, its results often require follow-up refinement by humans. For example, in the case study of stock price analysis, after generating a data video with an animated line chart of multiple companies and corresponding narration, the user may issue a follow-up instruction: “Emphasize Amazon’s data.” How might the system designer extend the existing multi-agent workflow to accommodate this new requirement?
@@
This design process yields a paradigm that aligns with P6, AI-driven Prompt Decomposition. By explicitly modeling entities and relations, the IAI framework helps system designers localize where and how human input should occur, and how to derive paradigms to support iterative communication between humans and AI. More broadly, this case demonstrates how the IAI model can serve as a reasoning tool for emerging scenarios: starting with entities, constraining relations along the two axes of interaction timing and user resources (Table 2), and applying cross-paradigm insights (Sec. 4.3) to guide concrete design choices.
Figure 8. Usage Scenario 3: Applying the IAI Model to Emerging Scenarios. For example, in a multi-agent system for animated data videos (Shen et al., 2024a), the IAI model guides the integration of user feedback for iterative refinement. This yields a paradigm aligned with P6 and demonstrates the IAI model can serve as a reasoning tool for emerging scenarios.
@@ 5.4. Usage Scenario 4: Deriving New Atomic Paradigm Graphs
Assume an individual enters a canteen and inspects the available dishes, the human–AI conversation is not initiated by the human through an explicit prompt (Fig. 9). Instead, an embodied AI assistant (e.g., embedded in AR glasses or a mobile app (Jiang et al., 2025; Wang et al., 2024b)) proactively perceives the environment (Sec. 4.3-C2), cross-references it with the individual’s dietary history, and initiates the interaction: “Hi, what do you want to eat today? Meat or vegetables?” Alongside this query, the system also generates an interactive widget (e.g., a slider) for specifying a preferred proportion of vegetables versus meat. Suppose the user selects 60% vegetables and 40% meat, and adds: “For meat, I’d like to try chicken, beef, or fish.” The AI then integrates this input with the detected canteen offerings to recommend a personalized list of top dishes. This new graph and usage scenario foregrounds AI-initiated, context-aware interaction, expanding the design space toward more proactive and situated human-GenAI collaborations (Sec. 4.3-C4), and more broadly, opening up fundamentally new paradigms of communication.
@@ 6.1. Reflection
What can our model contribute to interaction-augmented instructions? The IAI model contributes to both engineering practice and HCI theory-building. For practitioners, paradigm graphs serve as a design language that makes implicit design trade-offs explicit. Developers can compare alternatives, identify unexplored regions of the design space, and refine existing interfaces by recombining atomic paradigms. This resonates with prior calls to move beyond ad-hoc demonstrations toward systematic design frameworks in HCI (Beaudouin-Lafon, 2004). For researchers, our model complements prior taxonomies (Gao et al., 2024; Hu et al., 2025) and principle-based approaches (Riche et al., 2025) by introducing a formal representational structure that captures both entities and relations in human–GenAI workflows. Such formalization enables cumulative comparison across systems and provides a substrate for analytical and generative methods, similar to how prior models advanced earlier eras of HCI (Card et al., 1986; Beaudouin-Lafon, 2000; Card et al., 1999).

### v2

@@ 4.3. Cross-paradigm Insights and Suggestions
The model’s information-flow structure explains why interaction timing and artifact availability form the principal axes that differentiate the twelve paradigms; the paradigms themselves illustrate how recurring graph topologies map to distinct design affordances. In addition, our model disambiguates interface cases that prior frameworks conflate. The motivating examples in Fig. 2 (b–f) map cleanly to paradigm graphs P1, P3, P4, P7, and P9, respectively, as discussed above, demonstrating meaningful descriptive and discriminative power of the model. Building on this analytical foundation, we derive four concrete, model-grounded considerations that help designers reason about which paradigm class aligns with a given design situation based on interaction timing (C1) and artifact availability (C2), which specific paradigm within that class best matches user intent expression needs, as shown in Tab. 3 (C3), and how the IAI model and derived paradigms can be used generatively to propose new interfaces (C4).
@@
C1. Use interaction timing (interaction before/after invoking GenAI) to choose a paradigm class. Interaction timing indexes whether a task’s intent is knowable up front (pre-invocation) or emerges through exploration (post-invocation). Pre-invocation paradigms assume the user can articulate explicit and clear intent in advance for generating or manipulating artifacts   (Tab. 3 P1-P4), such as clear coding structure and logic (Feng et al., 2024), or explicit image editing areas and tasks  (Liu et al., 2025). Post-invocation paradigms assume underspecified or exploratory goals   (Tab. 3 P5-P12): the system first initiates interactions to further clarify user intents (Wang et al., 2024d) or produces outputs for steering and refinement (Zhang et al., 2025b).   Designers should align timing with the task’s nature based on the model’s discriminative power: employ pre-invocation paradigms for tasks requiring precision and auditable outcomes, but prioritize lightweight post-invocation paradigms when the goal is creative ideation and iterative refinement of an ambiguous intent.
@@
C2. Use artifact availability (prompt-only vs. artifact-grounded) to select how intent should be expressed. Instructions in natural language alone can be ambiguous (e.g., make the flower brighter” without specifying which flower). The presence of an artifact shifts intent expression from purely linguistic descriptions to concrete, grounded references   (Tab. 3 P4, P9-P12). By enabling users to select, highlight, or annotate artifact fragments, systems can reduce ambiguity and provide a stable foundation for commands (Singh et al., 2024; Chen et al., 2023). The consideration also aligns with design principles for AI-instruments by Riche et al.  (Riche et al., 2025), where they proposed that text prompts should be grounded in other artifacts.   Designers should identify whether their system will receive an initial artifact; if so, artifact-grounded paradigms offer lower ambiguity, richer referential precision, and reduced prompting burden.
@@
C3. Within a chosen class, refine toward a specific paradigm based on how intent should be encoded. In these design paradigms, a key design choice is how user intent is materialized to augment the original text prompt. The first way is to directly edit or extend the text prompt with interactions (Tab. 2-P1 and P5). An example is DreamSheets (Almeda et al., 2024), where users can combine prompt fragments easily with interactions. It provides a simple and direct way to improve original prompts but lacks the power to cater complex intent, such as non-linguistic or referential intent. To handle more complex intent, an approach is to leverage interactions to introduce additional non-linguistic information to text and form augmented instructions, such as structural or parametric information (Tab. 2-P2, P3, P6, P7, P9 and P11). For example, CoPrompt (Feng et al., 2024) allows users to drag and drop prompt fragments to form a multi-level list of instructions. It materializes the non-linguistic intent expression in a convenient way with suitable interaction design. Regarding the referential intent, the key interaction is to link text prompts with artifacts and generate augmented instructions (Tab. 2-P4, P10). It addresses the challenges in describing the linkage between texts and artifacts by materializing users’ intent with direct interactions on artifacts. Lastly, a unique case is to generate new interaction widget by text prompts and facilitate intent expression through direct manipulation (Tab. 2-P8, P12). GenAI creates reusable and persistent interactive widgets to materialize the meta intent and allows follow-up similar intent expression with simple interactions with widgets. A notable example is DynaVis (Vaithilingam et al., 2024). The meta intent like changing visual element colors can be materialized as a widget. The specific color to use can be directly selected by interactions.   The summarized design cheat sheet mapping common design situations to recommended paradigms is shown in Tab. 3. These differences correspond to distinct subgraph configurations in the atomic paradigm graphs. Designers can therefore match their intent-expression needs (e.g., decomposition, parameterization, selection specificity) to the paradigm whose graph topology supports that form of control.
Table 4.   Operationalizing Generative Power: Potential methods for designers and researchers to apply the IAI model and derived paradigms. In this table, starting points refer to whether the designer or researcher should look into IAI model or our summarized paradigms at the beginning of the usage method.
Design Goal Entry Point Core Workflow
Extend the Pipeline of an Existing Tool Paradigm 1. Identify the atomic paradigm of the existing tool, analyze the paradigm graph, and identify limitations.
(refer to Sec. 5.1) 2. When noticing the needs for an additional invocation of GenAI (e.g., asking additional questions for intent clarification), select and chain a complementary paradigm.
3. Verify the chained paradigm and introduce new designs.
Refine the Interaction Design of an Existing Tool Paradigm 1. Identify the atomic paradigm of the existing tool, analyze the paradigm graph, and identify limitations.
(refer to Sec. 5.2) 2. When it is not necessary for an additional invocation of GenAI (e.g., adding artifacts as part of intent for accurate intent expression), add or adjust targeted relations.
3. Reinterpret the new graph and introduce new designs.
Design a New Tool for an Emerging Scenario Model 1. Identify IAI model entities in the new scenario by considering scenario requirements and user resources.
(refer to Sec. 5.3) 2. Derive the relations based on interaction logic about timing and user resources.
3. Map to existing paradigms, get inspiration from example tools in other scenarios, and design the new tool.
Hypothesize and Research Novel Paradigms Paradigm/Model 1. Select base paradigms from existing ones or design base paradigms from scratch with the IAI model.
(refer to Sec. 5.4) 2. Experiment inverting, changing, and remixing different paradigms to form a new one. Identify potential advantages and usage scenarios of the new paradigm.
3. Study the new paradigm with tool prototypes and example application scenarios.
@@
C4. Reusing, chaining, and innovating atomic paradigms for adapting interaction design to new scenarios. Our summarized paradigms and the considerations C1-C3 above provide a concise mapping from task requirements to interface affordances.   Designers can start from intent clarity and artifact availability to select appropriate paradigms for interface implementation. For example, they can use the pre-invocation, artifact-grounded paradigm   (Tab. 3 P5) for tasks that start from existing artifacts and demand precision. They can apply post-invocation, prompt-driven paradigms   (Tab. 3 P5-P8) for exploration goals. The should also consider the intent type as C3 mentions. Crucially, the twelve paradigms are not mutually exclusive: they can be flexibly combined and chained within a single application, enabling systems to shift fluidly between scaffolding, refinement, and repurposing workflows (Gmeiner et al., 2025; Wang et al., 2024d). Effective interfaces thus treat paradigms as composable building blocks rather than rigid templates, supporting diverse user needs. For example, IntentTagger (Gmeiner et al., 2025) introduces small, atomic intent tags enabling micro-prompting and region-level edits, illustrating chaining and covering multiple paradigms (e.g., P2, P4, P5, P7, P9). Reusing and chaining paradigms are not the end. Comparing the summarized paradigms and the entire IAI model, we can notice there are plenty of potential atomic paradigms that have not been explored. In the future, interaction designers may start from existing paradigms to design new ones for innovation, indicating the model’s generative power.
@@ 5. Usage Scenario
The generative power of our proposed interaction-augmented instruction model lies in guiding both interaction paradigm and interactive tool innovation in human–GenAI communication (Beaudouin-Lafon et al., 2021). Together with our summarized existing paradigms, it can guide the iterative improvement of tools, facilitate the creation of novel interfaces tailored to emerging user scenarios, and enable generating new interaction paradigms beyond existing ones.   To illustrate the model’s generative power, we summarized four potential methods for designers and researchers to apply IAI models and summarized twelve paradigms for various needs, as shown in Tab. 4.
To demonstrate the four methods, we present four usage scenarios below to highlight how our model and paradigms can inspire, structure, and accelerate design decisions in real-world contexts. These scenarios were deliberately selected to cover the three fundamental ways in which our model can be used, forming a closed application space of design actions: (1) selecting and applying existing paradigms (Usage Scenarios 1 and 3), (2) refining existing paradigms by adjusting their graph structures (Scenario 2), and (3) generating new paradigms when existing ones are insufficient for a novel task (Scenario 4). The choice of application domains (e.g., data science, creativity, and everyday tasks for general users) is intentionally diverse to demonstrate that the IAI model is not domain-bound. Rather than exhaustiveness, our goal is to illustrate breadth and accessibility, enabling readers from different communities to understand how the model applies across contexts.
@@ 5.1. Usage Scenario 1: Extending Pipelines through Chained Paradigm Graphs
One way our model supports innovation is by extending existing human-GenAI interfaces through the chaining of additional atomic paradigm graphs onto their current workflows (Sec. 4.3-C4), with additional invocations of GenAI. This perspective treats current system interfaces not as static endpoints but as expandable foundations, where other paradigms can be strategically layered to address ambiguity and improve alignment
For example, when designing an ideal tool for data analysts to generate and edit visualizations, natural language can ease the burden of translating design requirements into initial visualizations, while graphical interactions can make subsequent adjustments and fine-grained editing more efficient. DynaVis (Vaithilingam et al., 2024) (Fig. 5-P8) illustrates this synergy by augmenting natural language interfaces with dynamic widgets that support iterative refinement. Yet, at the same time, early-stage ambiguity in user intent often remains a challenge. Fig. 7 illustrates this, imagine a data analyst at a movie company exploring market trends for the second quarter across U.S. cinemas. The analyst might ask: “Show me a trend of the most popular movie in the most recent three months.” Directly generating data insights to such fuzzy user questions (e.g., the most popular”) may not fully align with the user’s intent, frequently leading to time-consuming post-generation refinements. Using the IAI model and the atomic paradigm graphs, system designers can find and introduce a pre-generation disambiguation paradigm (P5, AI-driven Prompt Suggestion) before P8. Instead of immediately generating a visualization, the system could first initiate a clarification step, asking the analyst what “popular” should mean in this context, by ticket sales, IMDB rating, box office revenue, or number of screenings. This step structures the workflow to clarify ambiguous prompts upfront, improving alignment, reducing back-and-forth iterations, and making subsequent widget-driven refinement more precise.
@@
Figure 9. Usage Scenario 3: Applying the IAI Model to Emerging Scenarios. For example, in a multi-agent system for animated data videos (Shen et al., 2024b), the IAI model guides the integration of user feedback for iterative refinement. This yields a paradigm aligned with P6 and demonstrates the IAI model can serve as a reasoning tool for emerging scenarios.
@@ 5.2. Usage Scenario 2: Refining Paradigms by Adjusting Graph Structures
In addition to strengthening workflows at a macro level by chaining different paradigms, the flexibility of the relations among entities in the IAI model also enables micro-level edits within a single paradigm, without additional invocations of GenAI. Such refinements can be realized by adjusting the relations among entities in existing paradigms to improve precision and control.
@@ 5.3. Usage Scenario 3: Applying the IAI model to Emerging Scenarios
Consider Data Director (Shen et al., 2024b), a multi-agent system that automatically generates animated data videos from data tables with agent roles such as Data Analyst and Designer (Fig. 9). Although such end-to-end automation can rapidly deliver initial outputs, its results often require follow-up refinement by humans. For example, in the case study of stock price analysis, after generating a data video with an animated line chart of multiple companies and corresponding narration, the user may issue a follow-up instruction: “Emphasize Amazon’s data.” How might the system designer extend the existing multi-agent workflow to accommodate this new requirement?
@@
This design process yields a paradigm that aligns with P6, AI-driven Prompt Decomposition. By explicitly modeling entities and relations, the IAI framework helps system designers localize where and how human input should occur, and how to derive paradigms to support iterative communication between humans and AI. More broadly, this case demonstrates how the IAI model can serve as a reasoning tool for emerging scenarios: starting with entities, constraining relations along the two axes of interaction timing and user resources (Tab. 2), and applying cross-paradigm insights (Sec. 4.3) to guide concrete design choices.
Notably, while this scenario illustrates how the IAI model can support multi-agent workflow design, multi-agent systems are only one application context. The model does not formalize communication among multiple AI agents or include multiple G nodes within a single atomic paradigm. Each paradigm intentionally centers a single G node, and multi-agent workflows are expressed by chaining atomic paradigms if G is called multiple times for different purposes. Extending the model to represent shared or distributed Aug across agents is a promising future direction.
@@ 5.4. Usage Scenario 4: Deriving New Atomic Paradigm Graphs
Assume an individual enters a canteen and inspects the available dishes, the human–AI conversation is not initiated by the human through an explicit prompt (Fig. 10). Instead, an embodied AI assistant (e.g., embedded in AR glasses or a mobile app (Jiang et al., 2025; Wang et al., 2024c)) proactively perceives the environment (Sec. 4.3-C2), cross-references it with the individual’s dietary history, and initiates the interaction: “Hi, what do you want to eat today? Meat or vegetables?” Alongside this query, the system also generates an interactive widget (e.g., a slider) for specifying a preferred proportion of vegetables versus meat. Suppose the user selects 60% vegetables and 40% meat, and adds: “For meat, I’d like to try chicken, beef, or fish.” The AI then integrates this input with the detected canteen offerings to recommend a personalized list of top dishes. This new graph and usage scenario foregrounds AI-initiated, context-aware interaction, expanding the design space toward more proactive and situated human-GenAI collaborations (Sec. 4.3-C4), and more broadly, opening up fundamentally new paradigms of communication.
@@ 6.1. Reflection
What can our model contribute to interaction-augmented instructions? The IAI model contributes to both engineering practice and HCI theory-building.   It provides greater descriptive, discriminative, and generative power than methods in prior work, such as Gao et al.  (Gao et al., 2024) and Shen et al.  (Shen et al., 2025a) by introducing two innovations. First, the model explicitly encodes information flow in the paradigm graph. Each edge denotes a precise link between entities, making differences in timing, source, and composition of instructions formally visible. Another key contribution is treating Aug as the explicit instruction to GenAI for the expressiveness of the model. Without it, designs that combine text prompts, interactions, and artifacts in different ways collapse into the broad “non-text interactions” category. It cannot systematically distinguish important differences, such as that between P1 and P2. With Aug, these differences become structurally visible, varying how entities feed into GenAI, and in what order, systematically yielding distinct paradigms. This mechanism directly enables both our comparative taxonomy and its generative extension, where new interfaces emerge from principled recombinations of information flow patterns.
For practitioners, paradigm graphs serve as a design language that makes implicit design trade-offs explicit. Developers can compare alternatives, identify unexplored regions of the design space, and refine existing interfaces by recombining atomic paradigms. This resonates with prior calls to move beyond ad-hoc demonstrations toward systematic design frameworks in HCI (Beaudouin-Lafon, 2004). For researchers, our model complements prior taxonomies (Gao et al., 2024; Hu et al., 2025) and principle-based approaches (Riche et al., 2025) by introducing a formal representational structure that captures both entities and relations in human–GenAI workflows. For example, the three principles proposed in AI-Instruments (Riche et al., 2025) can be reinterpreted in IAI as properties of interaction flows rather than standalone design goals. Reification corresponds to whether and how intent is externalized into explicit intermediate representations (most notably Aug) that persist beyond a single interaction and can be inspected, reused, or chained. Reflection maps to interaction flows in which multiple alternatives are explicitly surfaced and negotiated over time, often through post-invocation paradigms where model outputs serve as anchors for iterative clarification. Grounding is realized through paradigms that bind instructions to concrete scopes (e.g., selected artifacts, regions, or prior outputs), thereby constraining interpretation through reference rather than description. Rather than treating these principles as orthogonal design heuristics, IAI shows how they manifest repeatedly across different atomic paradigms depending on interaction timing, available resources, and control flow. This shift reframes human–GenAI interaction from early-stage guideline analysis toward the concrete design and enactment of interaction structures, thereby enabling both actionable system designs and principled exploration of new paradigms. Such unified formalization enables cumulative comparison across systems and provides a substrate for analytical and generative methods, similar to how prior models advanced earlier eras of HCI (Card et al., 1986; Beaudouin-Lafon, 2000; Card et al., 1999).